Assume each firms long-run average cost are given by AC=10+0.002Q where Q is market output, so long-run costs are incre...

  1. Assume each firms long-run average cost are given by AC=10+0.002Qj45UAHDaaCRy0vSGOIxf98J8zY0I3pPcL09wJcDe where Q is market output, so long-run costs are increasing in market output. Market demand is given by QD=DP=1,050-50Pwydb1oOwMAXyWgblJU2tXsAAAAASUVORK5CYII=
    1. What is the long-run equilibrium price and market quantity?
    2. If market demand increases to QD=DP=1,600-50PAcy8y4zil4I4gAAAABJRU5ErkJggg== find the new long-run equilibrium price and market quantity.
    3. Graph these equilibrium outcomes and calculate the change in producer surplus between (a) and (b)
    4. If a tax of $5.50 per unit output is introduced find the price producers pay, price consumers pay and the new level of market output.
    5. What is the total tax revenue from the new tax? Assuming the market was in the equilibrium found in part (a) before the tax was introduced, what is the consumer and producers tax burden? What does the relative size of the producer’s and consumer’s tax burden tell you about the relative size of the elasticity of supply and elasticity of demand?
    6. Calculate the dead weight loss associated with the new tax.
Answer #1

Date long run Market anerage cost , AC = 10 + 0.002 a demand, Q = (011) = 1050-50P (a.) At long run equilibrium, we must = 7(6) 4 market demand increases do ao(P) = 1600 - 50P to new long run equilibrium 500 - 5000 = 1600 -sop. 550 P = 6600 p* = 12Im the ginen diagram pro en diagram surplus. producer area below the price line and above the supply curue. Howener in the r